# Regional Tiny Stories: Using Small Models to Compare Language Learning and Tokenizer Performance

### v1

@@ Abstract
Small Language Models (SLMs) offer efficient alternatives to LLMs for specific domains. The 2023 TinyStories study developed an English dataset that allows SLMs with 1–10 million parameters to produce coherent outputs. Our research expands this framework by translating the original dataset into Indian languages and creating synthetic data using LLMs. We focus on Hindi, Marathi, and Bengali, evaluating SLMs for regional language processing and understanding linguistic complexity. We show that SLMs efficiently process regional languages with significantly fewer parameters than LLMs, providing a complementary framework for “inference based evaluation” of tokenization strategies and linguistic complexity. Our analysis shows that language-specific tokenizers outperform general-purpose ones for Indian languages. Empirical validations, supported by information-theoretic and morphological analyses, provides fundamental understanding behind the better performance of Hindi models over Marathi and Bengali. Additionally, we show that synthetic datasets outperform translated content for training SLMs. Correlation analyses reveal cross-linguistic patterns and language-specific relationships between creativity, grammatical precision, and narrative completeness. These findings advance both the practical application of SLMs to underserved languages and our theoretical understanding of neural language development.
@@ 1 Introduction
  • Finally, we report and release our training data of $\approx$ 10M synthetic and translated stories in three Indian langauges for the broader community.
@@ 2.1 Training data preparation
Our research extends the TinyStories framework (Eldan & Li, 2023a) to explore simple, constrained narratives in multiple Indian languages through a two-phase approach: translating the original English TinyStories dataset (Eldan & Li, 2023b) into Indian languages, followed by generating additional synthetic data using LLMs while maintaining the original methodology’s constraints.
Figure 1: Schematic of model pipeline. (A) Dataset prepared through machine translation as well as generation using LLM, (B) Indic tokenizers are used to preprocess the Indian language stories, (C) A decoder only transformer architecture is used to train the model in each language, (D) Inference is evaluated by LLM on certain attributes
@@ 2.2 Training data evaluation
This pattern—high semantic similarity (BERTScore) with low lexical overlap (BLEU/METEOR)—indicates our dataset contains diverse lexical expressions of similar concepts. Such divergence occurs because morphologically rich Indian languages permit extensive variation in expressing equivalent meanings. The apparent metric ”anomaly” actually reveals a strength in our dataset: stories maintain semantic coherence while exhibiting rich linguistic variation, precisely the characteristics needed for robust language modeling. Rather than memorizing phrases, models learn to understand concepts expressed through diverse vocabulary and structures. This finding underscores both the challenge of evaluating Indian language generation and the benefit of our approach, which produces semantically coherent yet lexically diverse training examples that foster more generalizable language understanding.
@@ 3.1 Insights from our evaluation method
The impact of architectural scaling manifests distinctly across Hindi, Bengali, and Marathi languages. Hindi models achieve the strongest overall performance (8.164 at 1024/7 configuration), showing pronounced improvements in grammar (8.910) and fluency (8.580) while maintaining steady gains in context understanding with increased model size. Bengali models exhibit remarkably similar scaling characteristics (8.037 at 1024/7), suggesting fundamental similarities in structural complexity between these languages. In contrast, Marathi displays distinct scaling behavior, requiring larger architectures for comparable performance (7.881), suggesting its linguistic features demand additional model capacity for effective processing.
Examination of evaluation metrics reveals consistent improvement patterns in model evaluation loss across all languages, decreasing from approximately 1.4 to 0.5 in Hindi, indicating enhanced optimization in larger models. This trend appears most pronounced in Hindi and Bengali.
@@
  • Context understanding capabilities vary significantly: Hindi and Bengali demonstrate robust improvements with scale (reaching approximately 7.7), while Marathi exhibits more modest gains, pointing to distinct challenges in capturing contextual relationships.
  • Grammatical competence shows the most dramatic scaling improvements across all languages, though from different baselines – Hindi and Bengali achieve strong performance (approximately 8.4) even at modest scales, while Marathi requires larger models to reach comparable accuracy.
  • Optimal parameter allocation differs by language: Hindi models perform better with wider architecture (larger embedding dimensions), while Bengali models prefer a balanced approach with moderate width and depth.
@@
The hierarchical emergence of capabilities suggests that neural story generation follows a structured developmental pattern, similar to language acquisition, where simpler grammatical competencies precede more complex narrative abilities. This view aligns with the idea that language is acquired primarily through social interactions and pattern recognition in speech, rather than being driven by an innate, specialized ”language module.” Language learning emerges from general cognitive skills, such as intention-reading and pattern-finding, which underlie both grammatical and narrative development (Tomasello, 2003).
@@ 3.2 Comparing inference results of our Regional TinyStories with LLMs
In figures 2 and 3, we show the stories generated from our TinyStories 5M model and that of GPT-4o for Hindi language for the same prompt. The prompt which we gave in Hindi effectively translates to:
@@
We then ask GPT-4o to qualitatively compare between the 2 stories. Here is the output provided:
@@
3.3 Comparison between regional tokenizers
@@
3.4 Inference comparison for models trained on translated vs synthetic dataset
The 54M model with Sarvam tokenizer was chosen for inference evaluations on translated dataset and compared with synthetic. Check Doshi et al. More details are provided in Appendix F. May discuss other Tinystory papers. The 54M models with Sarvam tokenizer was chosen for inference evaluations on translated dataset and compared with those for the synthetic dataset. Check Doshi et al. More details are provided in Appendix F. The 54M models with Sarvam tokenizer was chosen for inference evaluations on translated dataset and compared with those for the synthetic dataset. Check Doshi et al. More details are provided in Appendix F. Lower inference scores for models trained on translated data could be explained using the following viewpoints:
  1. 1.
    Cultural biases: Source data culture transfers to target languages, including elements like foreign names that prevent models from generating culturally appropriate content (Holmström et al., 2023).
  2. 2.
    Grammatical and style issues: Languages express similar concepts with different strucutures and conventions, and often translations fail to adapt these nuances, producing unnatural text in the target language (Zhang & Toral, 2019).
  3. 3.
    Lastly, higher evaluation losses suggest difficulty in next token prediction which could stem from the noise introduced through the translation process, that complicates the task compared to working with original text (Boughorbel et al., 2024).
@@ 3.5 Tokenizer analysis with focus on language complexity
We employ a dual-perspective approach to quantitatively analyze the linguistic complexity of three major Indian languages: Hindi, Bengali, and Marathi. This analysis provides insights into both the intrinsic properties of these languages and the effectiveness of different tokenization strategies.
@@ 4 Conclusion
We demonstrated the TinyStories paradigm’s effectiveness for Indian languages, showing SLMs with just 5-50 million parameters can generate coherent outputs. Our largest model ($\approx 150M$ parameters) achieves 90 % of GPT-4o’s performance ($\approx 8$ vs $\approx 9$), despite being $10^{6}$ times smaller.
Regional TinyStories provides a novel framework for comparing language complexities and tokenizer efficiencies. We find that Hindi models using Sarvam perform the best. Its observed that Bengali favors creativity, Hindi has slightly weaker context-completeness links, and Marathi uniquely associates context with grammar. Through comprehensive evaluation, we established that Sarvam and SUTRA tokenizers outperform alternatives for Indian languages—the first study comparing regional tokenizers through inference evaluations rather than token fertility metrics. Our mechanistic analysis points toward higher entropy in tokenization behind the lower overall performance for Marathi compared to Hindi and Bengali, suggesting potential differences in how tokenizers deal with language complexity.
@@ Acknowledgement
We have released the entire code for dataset creation, model training and inference at:
(1) https://github.com/VizuaraAI/Tiny-Stories-Regional
(2) https://github.com/nirvan840/Vizuara-TinyStories-Regional
@@ B.0.2 Distributional Characteristics
The evaluation scores for all three languages exhibit distinct distributional patterns that reveal important aspects of model behavior (Figures 4-9). Below, we summarize key statistical properties observed across languages and metrics:
Table 8: Statistical Properties of Evaluation Metrics Across Languages
@@ B.1.1 Parameter Efficiency Across Languages
    Optimal Parameter Allocation: The inflection point in the performance-parameter curve occurs consistently around 54M parameters across all three languages, with minimal improvements beyond this threshold. However, the specific distribution of these parameters (between embedding dimension and layer depth) that yields optimal performance varies by language. Bengali achieves optimal performance with a balanced 512/6 configuration, while Hindi benefits more from increased width (1024/6) than depth.

### v2

@@ Abstract
The 2023 TinyStories study developed an English dataset that allows Small Language Models (SLMs) with 1–10 million parameters to produce coherent outputs matching those of LLMs. Our research expands this framework by creating translated as well as synthetically generated datasets in Indian languages. Using this new dataset , we demonstrate that SLMs efficiently process regional languages with significantly fewer parameters than LLMs, and additionally offer a complementary framework for “inference-based evaluation” of tokenization strategies and linguistic complexity. Our analysis reveals that language-specific tokenizers outperform general-purpose ones for Indian languages. Empirical validations, supported by information-theoretic and morphological analyses, provide insights into the superior performance of Hindi models over Marathi and Bengali. The study uncovers distinct cross-linguistic patterns: Bengali emphasizes creativity, Hindi excels in context understanding and grammar with model scaling, and Marathi requires larger models to capture its unique linguistic features. Optimal parameter allocation varies, with Hindi benefiting more from wider architectures and Bengali favoring a balanced approach. We also show that quality synthetic datasets outperform translated content for training SLMs by 15-30 % . These findings advance both the practical application of SLMs to underserved languages and our theoretical understanding of neural language development.
@@ 1 Introduction
  • Finally, we report and release our training data of $\approx$ 10M synthetic and translated stories in three Indian langauges along with the trained models for the broader community.
@@ 2.1 Training data preparation
Our research extends the TinyStories framework (Eldan & Li, 2023a) to explore simple, constrained narratives in multiple Indian languages through a two-phase approach: translating the original English TinyStories dataset (Eldan & Li, 2023b) into Indian languages, followed by generating additional synthetic data using LLMs while maintaining the original methodology’s constraints (Fig. 1).
Figure 1: Schematic of model pipeline. (A) Dataset prepared through machine translation as well as generation using LLM, (B) Indic tokenizers are used to preprocess the Indian language stories, (C) A decoder only transformer architecture (nanoGPT) is used to train the model in each language ; Inference is evaluated by LLM on linguistic attributes
@@ 2.2 Training data evaluation
This pattern—high semantic similarity (BERTScore) with low lexical overlap (BLEU/METEOR)—indicates our dataset contains diverse lexical expressions of similar concepts. Such divergence occurs because morphologically rich Indian languages permit extensive variation in expressing equivalent meanings. The apparent metric “anomaly” actually reveals a strength in our dataset: stories maintain semantic coherence while exhibiting rich linguistic variation, precisely the characteristics needed for robust language modeling. Rather than memorizing phrases, models learn to understand concepts expressed through diverse vocabulary and structures. This finding underscores both the challenge of evaluating Indian language generation and the benefit of our approach, which produces semantically coherent yet lexically diverse training examples that foster more generalizable language understanding.
@@ 3.1 Insights from our evaluation method
Model evaluation loss consistently decreases across all languages, with Hindi dropping from about 1.4 to 0.5, indicating improved optimization in larger models. This trend is most pronounced in Hindi and Bengali, suggesting better learning. Architectural scaling impacts Hindi, Bengali, and Marathi differently. Hindi models achieve the best performance (8.164 at 1024/7 configuration) with notable improvements in grammar (8.910) and fluency (8.580), and steady gains in overall score as model size increases. Bengali models show similar scaling characteristics (8.037 at 1024/7), hinting at structural similarities between these languages and indeed they belong to fusional languages. Conversely, Marathi which is an agglutinative language, requires still larger models for similar performance, indicating its linguistic features need more model capacity for effective processing.
@@
  • Context understanding capabilities vary significantly: Hindi and Bengali demonstrate robust improvements with scale (reaching approximately 7.7 at 1024/7), while Marathi exhibits more modest gains, pointing to distinct challenges in capturing contextual relationships.
  • Grammatical competence shows the most dramatic scaling improvements across all languages, though from different baselines – Hindi and Bengali achieve strong performance (approximately 8.4 at 64/6) even at modest scales, while Marathi requires larger models (512/2) to reach comparable accuracy.
  • Optimal parameter allocation differs by language: Hindi models keep gaining with wider architecture (more embeddings), while Bengali models prefer a balanced approach with moderate width and depth.
  • Interestingly, increasing embedding dimensions beyond 512, largely stymies the percentage improvement in attributes like context, grammar and overall score. This is most evident in Marathi and Bengali but further increase in layer depth allows some improvements.
@@
The hierarchical emergence of capabilities suggests that neural story generation follows a structured developmental pattern, similar to language acquisition, where simpler grammatical competencies precede more complex narrative abilities. This view aligns with the idea that language is acquired primarily through social interactions and pattern recognition in speech, rather than being driven by an innate, specialized “language module.” Language learning emerges from general cognitive skills, such as intention-reading and pattern-finding, which underlie both grammatical and narrative development (Tomasello, 2003).
@@ 3.2 Comparing inference results of our Regional TinyStories with LLMs
In Figs. 2 & 3, we show the stories generated from our TinyStories 5M model and that of GPT-4o for Hindi language for the same prompt. The prompt which we gave in Hindi effectively translates to:
@@
We then ask GPT-4o to qualitatively compare between the 2 stories. Here is the output :
@@
3.3 Regional tokenizers perform better
@@
3.4 Models trained on translated data have lower inference evaluations
The 54M model with Sarvam tokenizer was chosen for inference evaluations on translated datasets and compared with those trained on synthetic data. We observe that these models have higher evaluation loss for same training duration and have poorer inference evaluation scores , suggesting that translated data quality is lower compared to synthetic. For e.g., the Hindi model has an overall score of 6.39 against 8.16 observed for model trained with synthetic data (Appendix F). Reduced performance, upon using translated data, was also noticed in one previous report (Boughorbel et al., 2024). This could be explained using the following viewpoints: (1) Cultural biases: Source data culture transfers to target languages, including elements like foreign names that prevent models from generating culturally appropriate content (Holmström et al., 2023). (2) Grammatical and style issues: Languages express similar concepts with different strucutures and conventions, and often translations fail to adapt these nuances, producing unnatural text in the target language (Zhang & Toral, 2019). (3) Lastly, higher evaluation losses suggest difficulty in next token prediction which could stem from the noise introduced through the translation process, that complicates the task compared to working with original text (Boughorbel et al., 2024).
@@ 3.5 Tokenizer analysis with focus on language complexity
We use a dual-perspective approach to quantify the linguistic complexity of Hindi, Bengali, and Marathi, revealing their intrinsic features for evaluating tokenization strategies.
@@ 4 Conclusion
We demonstrated the TinyStories paradigm’s effectiveness for Indian languages, showing SLMs with just 5-50 million parameters can generate coherent outputs. Our largest model ($\approx 150M$ parameters) achieves 90 % of GPT-4o’s performance ($\approx 8$ vs $\approx 9$ on overall evaluations), despite being $10^{6}$ times smaller.
Further Regional TinyStories provides a novel approach for comparing language complexities and tokenizer efficiencies based on our inference evaluation framework. We find that Hindi models using Sarvam perform the best. Its observed that Bengali favors creativity, Hindi has slightly weaker context-completeness links, and Marathi uniquely associates context with grammar, suggesting distinct learning capabilities. Through comprehensive evaluation, we established that Sarvam and SUTRA tokenizers outperform alternatives like Tiktoken —the first study comparing regional tokenizers through inference evaluations rather than token fertility metrics. Our mechanistic analysis points toward higher entropy in tokenization behind the lower overall performance for Marathi compared to Hindi and Bengali, suggesting potential differences in how tokenizers deal with language complexity.
@@ Acknowledgement
We publicly release all our codes and datasets used for the experiments described in this work as well as selected trained models, to allow for external validation and reproducibility :
@@ B.0.2 Distributional Characteristics
The evaluation scores for all three languages exhibit distinct distributional patterns that reveal important aspects of model behavior (Figures 4-9). Below, we summarize key statistical properties observed across languages and metrics.
Table 8: Statistical Properties of Evaluation Metrics Across Languages for 54M models
@@ B.1.1 Parameter Efficiency Across Languages
    Optimal Parameter Allocation: The inflection point in the performance-parameter curve occurs consistently around 54M parameters across all three languages, with minimal improvements beyond this threshold. However, the specific distribution of these parameters (between embedding dimension and layer depth) that yields optimal performance varies by language. Bengali achieves optimal performance with a balanced 512/6 configuration, while Hindi benefits more from increased width (1024/7) than depth.
@@
Figure 4: Percentage Improvement in Contextual Comprehension, Grammatical Accuracy, and Overall Performance from Baseline Scores across Hindi models detailed in Table 1
Figure 5: Percentage Improvement in Contextual Comprehension, Grammatical Accuracy, and Overall Performance from Baseline Scores across Marathi models detailed in Table 2
Figure 6: Percentage Improvement in Contextual Comprehension, Grammatical Accuracy, and Overall Performance from Baseline Scores across Bengali models detailed in Table 3
@@ Appendix F Details on translated dataset and inference evaluations
F.1 Translation method
The original TinyStories dataset in English (Eldan & Li, 2023b) has a train split that contains 2,119,719 stories, with 320,470 duplicates, reducing unique stories to 1,799,249 (15.12% duplicates). This may impact model training, potentially skewing results.
Examples of stories appearing three times include :
  • ”Lily and Ben were playing in the park. They saw a fox hiding in the bushes. The fox had red fur and …”
  • ”Tom is a boy who likes to play with his lab. His lab is a big, black dog who can do amazing tricks. …”
  • ”Sara loved her pet cat, Lily. Lily was soft and fluffy and liked to sleep on Sara’s bed. Sara liked …”
  • ”Sam was sick. He had a bad cough and a sore throat. He did not like being sick. He wanted to play wi…”
Validation split has 21,990 unique stories without duplicates, ensuring reliable evaluation. However, cross-split analysis shows 6,601 stories appear in both splits, which could inflate model performance metrics. Addressing these duplicates is crucial for accurate model comparison between translated and synthetic data, enhancing the robustness of our findings in regional language modeling. Hence, we removed the duplicates and merged the two datasets. Additionally, due to resource constraints we could only translate for Hindi and Bengali. Apart from the previously mentioned issue with duplicates, we observed some data loss in translation due to issues with the freely available Google Translate API, resulting in approximately 1.9 M unique stories in Hindi and Bengali. Due to severe resource constraint, we could not translate to Marathi.
It was decided to use GPT4o for rating the translation quality based on recent reports (Kocmi & Federmann, 2023; Jiao et al., 2023) of GPT4 matching or outperforming Google Translate in case of non-English languages. In the context of evaluating translation models for Hindi and Bengali, using GPT4o, several multilingual models were assessed to identify the most effective translation method. The models include mBART (Liu et al., 2020), IndicTrans2 (Gala et al., 2023), mT5 (Xue et al., 2020), Helsinki-NLP OPUS MT (Tiedemann et al., 2023), IndicBART (Dabre et al., 2021), NLLB (Team et al., 2022), and M2M100 (Fan et al., 2021). Each model offers distinct capabilities, particularly in handling translations between English and Indic languages. GPT4o was given original English text and asked to translate to the target Indic language. Following this, it was provided with one of the machine translations and asked to rate it out of 10, using the following prompt :
I have used a machine translation model to translate the original story. On a scale of 1-10 evaluate the translation quality of the story, with respect to your translation, 1 being very bad and 10 being of the same quality as yours. Remember that ”quality” here does not mean the same words but meaningfully retaining the same context and fluency. Also, point out each instance where there is a mistake in translation.
For Bengali translation, IndicTrans2 achieved an average score of 7/10 on 100 stories. Google Translate, although widely used, sometimes produced translations with contextual inaccuracies and verb errors. Still it was rated at an average score of 8 by GPT-4o. Hence it was chosen for Bengali. On the other hand, the NLLB model received an average score of 7/10. GPT-4o noted that while the NLLB translations were generally good, they lacked the fluency and natural phrasing found in Google Translate. Specific issues were identified with verb tenses and word choices, leading to minor awkwardness in sentence structures.
We used a Python library called DeepTranslate to perform Google Translate API calls. This ran on CPU hardware but we performed several optimizations to improve translation speed, namely batch processing of stories, rate limiting at 5 requests per second, and robust fault tolerance measures. These measures included automatic retries with exponential backoff, an LRU cache for successful translations, and regular intermediate saves. Quality control involved clear story demarcation, UTF-8 encoding, and progress tracking.
However, for Hindi translations both NLLB and Google Translate produced scores around 8.5. Hence, the total Hindi transaltion was split between these two, primarily due to resource planning regarding API calls for Google Translate and GPU time for NLLB.
Overall, these evaluations demonstrate the strengths and weaknesses of different models in translating between English and Indic languages, highlighting the importance of choosing a model that balances semantic accuracy with natural language fluency. This analysis is critical for understanding the challenges and opportunities in model translation quality, particularly in academic contexts where precise language use is essential.
F.2 Inference scores
Dataset used Eval Loss Context Completeness Creativity Fluency Grammar Overall
Hindi
Translated data 1.3849 6.178 5.806 5.938 6.597 7.438 6.391
Synthetic data 0.518 7.734 7.783 7.806 8.554 8.912 8.158
Bengali
Translated data 1.4939 6.879 6.598 6.462 7.339 8.122 7.080
Synthetic data 0.569 7.507 7.645 7.693 8.420 8.816 8.016
Table 14: Inference score evaluations for Hindi and Bengali in case of models trained with translated vs synthetic dataset. Model configuration : 6 layers, 8 attention heads, 512 hidden embeddings. 3000 stories were evaluated for each model. Its observed that models trained with translated data have reduced performance.
Boughorbel et al. (2024) explored how machine-translated data can be used to train language models for generating stories in Arabic, and pinpointed problems related to linguistic and cultural biases present in the translated data. They suggested additional pre-training of the models using a limited amount of high-quality synthetic data and examined the impact of this approach using Dictionary Learning techniques. They further reported that the resulting Sparse Auto-Encoders demonstrated a change in the features they learn, showing improved linguistic characteristics and less cultural bias. In any event, we have shown that training on fully synthetic data that has been generated with high quality vetting process can indeed result in significant performance improvements as shared in Table 14. The drop in quality of inference for Hindi compared to Bengali could be due to two separate translation means employed for generating Hindi translated data. We do not have the resource at present to thoroughly evaluate this.